OpenStudy (anonymous):

Find the area of the surface obtained by rotating the curve of parametric equations: x=3t−3/3(t^3) y=3t^2 0

OpenStudy (anonymous):

So I set up my integral like so : \[2\pi \int\limits_{0}^{1}(3t-t^3)\sqrt{(-3t^2+3)^2+(6t)^2}\]

OpenStudy (anonymous):

After evaluating it I get an answer of 11pi but that is incorrect. Any help would be really appreciated!

OpenStudy (accessdenied):

i agree with most of it: what was your reasoning for the choice of using x=3t - t^3 ds ?

OpenStudy (anonymous):

Because it's being rotated about the x-axis

OpenStudy (accessdenied):

hm.. but if we revolve it around the x-axis:

OpenStudy (accessdenied):

 this radius is the y-value, correct?

OpenStudy (anonymous):

Yes

OpenStudy (accessdenied):

basically, the surface area is being defined as the sum of each circumference of a circle: 2 pi r which we add up through all small lengths ds along the curve (which we convert to dt) we just said that radius was y? integral of 2 pi * y * ds rather than 2 pi x

OpenStudy (accessdenied):

so i think if you use y = 3t^2 rather than x = 3t - t^3, you should get the correct integral and correct answer (looking at http://tutorial.math.lamar.edu/Classes/CalcII/ParaSurfaceArea.aspx it is also the given formula for rotation about x-axis)

OpenStudy (anonymous):

Ohhh okay I see. I got the right answer now! Thanks so much for the help.

OpenStudy (accessdenied):

glad to help! :)